# Forward Ray Tracing and Hot Spots in Kerr Spacetime

### v1

@@ Abstract
Hotspots, often characterized as point-like emissions, frequently manifest near black holes, displaying significantly increased luminosity compared to the surrounding accretion flow. It is noteworthy that these hotspots regularly occur outside the black hole at the center of the Milky Way. Light rays emitted from these sources follow complex trajectories around the black hole, ultimately arriving at distinct locations on the observer’s image plane. To extract precise spacetime information, including the black hole mass, spin, and inclination angle, it is crucial to accurately resolve both the direct emission and its higher-order images, despite the latter’s intensity suppression. To enhance the accuracy and efficiency of modeling and analyzing these hotspots, we introduce a forward ray tracing method, a departure from the traditional backward ray tracing approach. By utilizing conserved quantities in Kerr spacetime, this method initiates geodesics from a specified emission point near the black hole and terminates them at a distant observer, effectively capturing multiple images. By introducing perturbations to these geodesics, we map finite-size emissions to distinct regions on the image plane, allowing for the quantification of image shapes and amplification rates. This approach facilitates efficient spacetime tomography and hotspot localization, leveraging observations from the Event Horizon Telescope and its next-generation upgrades.
@@ I Introduction
Hotspots, characterized by transient and localized emissions that significantly overshadow their surrounding areas, represent a promising source for detecting light echoes Broderick and Loeb (2005, 2006); Moriyama and Mineshige (2015); Saida (2017); Moriyama et al. (2019); Tiede et al. (2020); Wong (2021); Chesler et al. (2021); Levis et al. (2022). Their intense emissions and pronounced variability facilitate the clear distinction between both their direct emissions and the secondary echoes against the background of accretion flows or jets. In the case of Sgr A, flares from hotspots are observed on a daily basis and span the electromagnetic spectrum Trippe et al. (2007); Marrone et al. (2008); Witzel et al. (2021), including the millimeter band Doeleman et al. (2008); Fish et al. (2008); Doeleman et al. (2009); Johnson et al. (2014); Wielgus et al. (2022a, b), near-infrared Genzel et al. (2003); Eckart et al. (2006); Meyer et al. (2006); Zamaninasab et al. (2010); Do et al. (2019b); Abuter et al. (2023), and X-rays Baganoff et al. (2001); Porquet et al. (2003); Kusunose and Takahara (2011); Karssen et al. (2017); Haggard et al. (2019); Andrés et al. (2022). These hotspots can be situated remarkably close to the BH, within a few gravitational radii Gravity Collaboration et al. (2018); Bauböck et al. (2020); Jiménez-Rosales et al. (2020); Wielgus et al. (2022b); Abuter et al. (2023). Magnetic reconnection within the plasma is a plausible mechanism for their formation, as indicated by general relativistic magnetohydrodynamics (GRMHD) simulations Younsi and Wu (2015); Dexter et al. (2020); Ripperda et al. (2020); Porth et al. (2021); Dexter et al. (2020); Ripperda et al. (2020); Chatterjee et al. (2021); Ripperda et al. (2022); Vos et al. (2024); Xi and Feng (2024). Observations in astrometry, polarimetry, and light curves of the flares enable the reconstruction of hotspot motion, providing a unique opportunity to probe plasma dynamics and spacetime by exploring different regions around the BH Tiede et al. (2020); Bauböck et al. (2020); Cardoso et al. (2021); Jiménez-Rosales et al. (2020); Wielgus et al. (2022b); Vos et al. (2022); von Fellenberg et al. (2023); Levis et al. (2024); Abuter et al. (2023); Vincent et al. (2024); Kocherlakota et al. (2024); Antonopoulou and Nathanail (2024); Yfantis et al. (2024a).
@@
Extracting spacetime information from a hotspot, particularly through its echoes, which can unveil universal features independent of the plasma composition, necessitates simulating its emission within the accretion flow around the SMBH, as exemplified in Refs. Broderick and Loeb (2005); Meyer et al. (2006); Trippe et al. (2007); Hamaus et al. (2009); Zamaninasab et al. (2010); Tiede et al. (2020); Jiménez-Rosales et al. (2020); Matsumoto et al. (2020); Dokuchaev and Nazarova (2021); Bauböck et al. (2020); Levis et al. (2022); Wielgus et al. (2022b); Vos et al. (2022); Rosa et al. (2022); Guo et al. (2023); Aimar et al. (2023); Vincent et al. (2024); Rosa et al. (2023); Najafi-Ziyazi et al. (2023); Yfantis et al. (2024b); Levis et al. (2024); Chen et al. (2024a); von Fellenberg et al. (2023); Rosa et al. (2024); Chen et al. (2024b); Kocherlakota et al. (2024); Antonopoulou and Nathanail (2024); Yfantis et al. (2024a). This task predominantly employs backward ray tracing, a method that traces geodesics from the observer’s image plane back towards the BH Bardeen (1973); Luminet (1979), and is a foundational technique in most ray tracing and radiative transfer packages Dexter (2016); Moscibrodzka and Gammie (2018); Bronzwaer et al. (2018); Pihajoki et al. (2018); Younsi et al. (2019); Bronzwaer et al. (2020); White (2022); Aimar et al. (2024); Prather et al. (2023); Huang et al. (2024). However, efficiently simulating hotspot echoes presents challenges. The implementation of a finite grid on the image plane complicates the precise targeting of point-like sources. While increasing the emission volume may assist, the probability of capturing higher-order images exponentially decreases with each additional image order. Moreover, accurately reconstructing the spatial distribution of emissions within the hotspot is challenging, as only a limited number of geodesics intersect the emission volume. These challenges also extend to the precise determination of hotspot locations and BH parameters from observational data.
@@
In this study, we introduce a novel framework termed forward ray tracing, which initiates geodesics from a specified emission point near a BH and terminates them at a distant observer, effectively capturing multiple images. Utilizing conserved quantities within Kerr spacetime Carter (1968); Gralla and Lupsasca (2020b); Lupsasca et al. (2024), this approach sets the initial direction of the geodesics based on each half-orbit count, $N$. This method is highly efficient and streamlined, ideal for modeling emissions from point-like sources. Additionally, by integrating perturbation theory into forward ray tracing, we map a finite emission volume to specific areas on the image plane, allowing us to quantify the shapes and amplification rates of images. Our findings reveal that higher-order images undergo spatial compression perpendicular to the critical curve due to the photon ring’s instability, while in the parallel direction, these images may either compress or expand depending on the specific location of the emission. We apply our forward ray tracing method to realistic observation scenarios involving highly time-variable hotspots near M87 or Sgr A, conducting analyses to determine the hotspot location and the BH parameters. This approach demonstrates significant promise for elucidating the BH spin with the next-generation EHT (ngEHT) Emami et al. (2023); Johnson et al. (2023); Ayzenberg et al. (2023).
@@
The structure of this paper is as follows: Section II offers an overview of null geodesics within Kerr spacetime and the backward ray tracing method. Section III introduces the concept of forward ray tracing and its novel application in generating direct and higher-order images from a point source. In Section IV, a perturbative mapping method is presented, which translates emissions from a finite volume to an array of regions on the image plane, allowing for a systematic analysis of image distortion as the order increases. Section V explores the application of forward ray tracing to spacetime tomography and hotspot localization. Section VI summarizes our findings and discusses potential future directions for both theoretical and observational studies. Numerical code KerrP2P Ker (2024) used in this study is available  [^1] KerrP2P: https://github.com/AuroraDysis/KerrP2P. Throughout this study, we adopt a system of geometrized units where the gravitational constant $G$ and the speed of light $c$ are both set to $G=c=1$.
@@ II Null Geodesics in Kerr Spacetime and Backward Ray Tracing
This section provides an overview of null geodesics in Kerr spacetime, focusing on the integral formalism as outlined in Refs. Gralla and Lupsasca (2020b, a). We explore the trajectories of light rays that reach a distant observer, examining their motion and detailing how they are governed by two conserved quantities in Kerr spacetime. These quantities are then mapped to specific coordinates on the image plane. We conclude this section with a brief discussion of the traditional backward ray tracing method used to calculate BH images, paving the way for the introduction of the new forward ray tracing method in the following section.
@@ II.1 Null Geodesics in Kerr Spacetime
where the superscripts ${+,-,f,s}$ correspond to the evaluation of $\mathcal{G}_{j}$ at $\theta=\theta_{+}$, $\theta_{-}$, $\theta_{f}$, and $\theta_{s}$, respectively. Ref. Gralla and Lupsasca (2020b) provides the analytical expressions for $\mathcal{G}_{j}$.
@@
where $\gamma$ is the Lyapunov exponent characterizing the instability of bound orbits, as defined in Ref. Cardoso et al. (2009); Gralla and Lupsasca (2020a). Conserved quantities of critical photons correspond to a curve known as the “critical curve” on the $\lambda$-$\eta$ plane, which delineates two distinct regions in the positive-$\eta$ portion of the $\lambda$-$\eta$ plane, each corresponding to a different type of radial motion.
The classification of radial motion leverages the roots of the radial potential $\mathcal{R}(r)$, denoted as ${r_{1},r_{2},r_{3},r_{4}}$, arranged such that $\operatorname{Re}r_{1}\leq\operatorname{Re}r_{2}\leq\operatorname{Re}r_{3}\leq\operatorname{Re}r_{4}$. Detailed analytical expressions for these roots are available in Ref. Gralla and Lupsasca (2020b). The behavior of $r_{4}$ is crucial in determining the nature of radial motion. When $(\lambda,\eta)$ falls inside the critical curve, $r_{4}$ is either complex or real but less than the outer horizon at $r_{h}\equiv M+\sqrt{M^{2}-a^{2}}$. In this scenario, the radial potential $\mathcal{R}(r)$ does not admit zeros outside the horizon, indicating no radial turning points are present, and only photons with $\nu_{r}^{s}=+1$ can escape. We denote the number of radial turning points encountered by $w$, which is zero ($w=0$) in this scenario.
@@
Table 1: Classification of radial motion capable of escaping the BH, delineated by the number $w$ of radial turning points encountered by the photon. ‘CC’ denotes the critical curve.
@@
with the superscripts $f$, $s$, and $4$ indicating the evaluation of $\mathcal{I}_{i}$ at $r=r_{f}$, $r_{s}$, and $r_{4}$, respectively. The formulas for the antiderivative $\mathcal{I}_{i}(r)$ (for $i=r,\phi,t$) are detailed in Ref. Gralla and Lupsasca (2020b).
@@ II.2 Image Plane
In practice, simulating an image of a BH surrounded by an accretion flow often involves backward ray tracing. This method launches geodesics from a grid on the observer’s image plane, parameterized by $(\alpha,\beta)$, back towards the BH Bardeen (1973); Luminet (1979). The geodesics’ endpoints either fall into the BH or extend to a large Mino time for photons situated outside or on the critical curve. The next step involves calculating the emissivity and absorption coefficients along each geodesic. These values are then integrated back to the observer, factoring in both gravitational and Doppler redshift effects, in a procedure termed covariant radiative transfer Gammie and Leung (2012); Dexter (2016).
@@ III Forward Ray Tracing and Light Echoes
The method of backward ray tracing, while effective for various applications, falls short in efficiently simulating higher-order images emanating from point-like sources, such as hotspots. This inefficiency stems from the challenge of aligning light rays from a finite observer grid precisely with the emission point. Additionally, the photon ring’s inherent instability exacerbates this issue, as the likelihood of accurately capturing these light rays decreases exponentially with the increase in the order number $N$. To overcome these challenges and improve the simulation of images from hotspots, we introduce a new approach called forward ray tracing. This section outlines this innovative method and demonstrates its utility in compiling images from a point source.
@@ III.1 Forward Ray Tracing
This subsection elaborates on the methodology behind our forward ray tracing approach, which utilizes the integral formalism discussed in Refs. Gralla and Lupsasca (2020b, a) and previously mentioned in Sec. II.1. This approach, initiating at the source point, computes light rays that ultimately converge at the observer, yielding observable images on the plane described in Sec. II.2. Note that our discussion is confined to scenarios where $\eta>0$, thus omitting vortical motion.
@@ III.1.1 Boundary Conditions to Fix Light Rays
Within the refined parameter space, each $(\lambda,\eta,\nu_{r}^{s},\nu_{\theta}^{s})$ combination corresponds to a light ray directed towards a distant observer. Next, we consider the light ray’s final direction. We begin by setting the final radial position to $r_{f}=r_{o}$, which allows us to determine the Mino time $\tau\equiv I_{r}$ through Eq. (16). Subsequently, the final polar angle $\theta_{f}$ is obtained from Eq. (11). This enables us to calculate the number of angular turnings $m$ encountered during propagation
@@
The concluding requirement for a light ray is to satisfy the imaging conditions
@@
Figure 1: Re-parameterization of conserved quantities of geodesics used in root-finding. The example shown is for a BH with $a/M=0.8$. The blue curve represents the critical curve corresponding to bound orbits, with arrows indicating the direction of increasing $\tilde{r}$.
In practice, to efficiently identify the roots with high precision, we transform the variables $(\lambda,\eta)$ into a new pair of parameters. We begin by introducing $q\equiv\sqrt{\eta}$, ensuring $\lambda$ and $q$ share the same dimensionality. The value of $q$ along the critical curve is denoted as $\tilde{q}\equiv\sqrt{\tilde{\eta}}$. We then assess the deviation of $(\lambda,q)$ from their critical curve counterparts $(\tilde{\lambda},\tilde{q})$ via the parameterization
@@
represents the outward-pointing unit normal vector to the critical curve $(\tilde{\lambda},\tilde{q})$ on the $\lambda$-$q$ plane. The parameter $d$ indicates the distance of $(\lambda,q)$ from the critical curve, with regions outside and inside the curve distinguished by $d>0$ and $d<0$, respectively. This approach allows each point $(\lambda,q)$ to be uniquely mapped to a distance $d$ and the radius $\tilde{r}$ of the corresponding critical curve, as illustrated in Fig. 1. Given that higher-order images asymptotically approach the critical curve, we opt for $\log_{10}|d/M|$ as the variable to accommodate the asymptotic behavior of radial motion. The root-finding process is then conducted on the $\tilde{r}$-$\log_{10}|d/M|$ plane, with a total of eight possible combinations of sign variables $(\nu_{r}^{s},\nu_{\theta}^{s},\operatorname{sgn}(d))=(\pm 1,\pm 1,\pm 1)$, which should be solved one by one. For each combination of sign variables, the contours for $\theta_{f}=\theta_{o}$ and $\phi_{f}=\phi_{o}+2k\pi\ (k\in\mathbb{Z})$ are drawn, with their intersections denoting the images. Note that for a Schwarzschild BH, since $\tilde{r}$ is always $3M$, finding a root in the one-dimensional space of $\log_{10}|d/M|$ suffices. Equivalently, this can be done by restricting to the equatorial plane, setting $\eta=0$, and solving for the remaining conserved quantity $\lambda$. See Sec. IV.2 for details.

### v2

@@ Abstract
Hotspots, often characterized as pointlike emissions, frequently appear near black holes with significantly enhanced luminosity compared to the surrounding accretion flow. Notably, such hotspots are regularly observed near the black hole at the center of the Milky Way. Light rays emitted from these sources follow complex trajectories around the black hole before reaching distinct locations on the observer’s image plane. Precisely resolving both direct emissions and their higher-order images—despite the latter’s intensity suppression—is essential for extracting detailed spacetime information, including the black hole’s mass, spin, and inclination angle. To improve the accuracy and efficiency of hotspot modeling, we develop a forward ray tracing method based on the analytic integral solution of Kerr geodesics, leveraging conserved quantities. Our approach traces geodesics from a given emission point near the black hole to a distant observer, effectively capturing multiple images with a tailored parametrization scheme for root-finding. By perturbing these geodesics, we map finite-size emissions to distinct regions on the image plane, enabling the quantification of image shapes and amplification rates. This method not only enhances the identification of strongly lensed photons from black holes but also enables efficient spacetime tomography and hotspot localization, leveraging observations from the Event Horizon Telescope and its upcoming next-generation upgrades.
@@ I Introduction
Hotspots, characterized by transient and localized emissions that significantly overshadow their surrounding areas, represent a promising source for detecting light echoes Broderick and Loeb (2005, 2006); Moriyama and Mineshige (2015); Saida (2017); Moriyama et al. (2019); Tiede et al. (2020); Wong (2021); Chesler et al. (2021); Levis et al. (2022) compared to the extended accretion flow Cárdenas-Avendaño et al. (2024). Their intense emissions and pronounced variability facilitate the clear distinction between both their direct emissions and the secondary echoes against the background of accretion flows or jets. In the case of Sgr A, flares from hotspots are observed on a daily basis and span the electromagnetic spectrum Trippe et al. (2007); Marrone et al. (2008); Witzel et al. (2021), including the millimeter band Doeleman et al. (2008); Fish et al. (2008); Doeleman et al. (2009); Johnson et al. (2014); Wielgus et al. (2022a, b), near-infrared Genzel et al. (2003); Eckart et al. (2006); Meyer et al. (2006); Zamaninasab et al. (2010); Do et al. (2019b); Abuter et al. (2023), and x-rays Baganoff et al. (2001); Porquet et al. (2003); Kusunose and Takahara (2011); Karssen et al. (2017); Haggard et al. (2019); Andrés et al. (2022). These hotspots can be situated remarkably close to the BH, within a few gravitational radii Gravity Collaboration et al. (2018); Bauböck et al. (2020); Jiménez-Rosales et al. (2020); Wielgus et al. (2022b); Abuter et al. (2023). Magnetic reconnection within the plasma is a plausible mechanism for their formation, as indicated by general relativistic magnetohydrodynamics (GRMHD) simulations Younsi and Wu (2015); Dexter et al. (2020); Ripperda et al. (2020); Porth et al. (2021); Dexter et al. (2020); Ripperda et al. (2020); Chatterjee et al. (2021); Ripperda et al. (2022); Vos et al. (2024); Xi and Feng (2024). Observations in astrometry, polarimetry, and light curves of the flares enable the reconstruction of hotspot motion, providing a unique opportunity to probe plasma dynamics and spacetime by exploring different regions around the BH Tiede et al. (2020); Bauböck et al. (2020); Cardoso et al. (2021); Jiménez-Rosales et al. (2020); Wielgus et al. (2022b); Vos et al. (2022); von Fellenberg et al. (2023); Levis et al. (2024); Abuter et al. (2023); Vincent et al. (2024); Huang et al. (2024a); Kocherlakota et al. (2024); Antonopoulou and Nathanail (2024); Yfantis et al. (2024a).
@@
Extracting spacetime information from a hotspot, particularly through its echoes, which can unveil universal features independent of the plasma composition, necessitates simulating its emission within the accretion flow around the SMBH, as exemplified in Refs. Broderick and Loeb (2005); Meyer et al. (2006); Trippe et al. (2007); Hamaus et al. (2009); Zamaninasab et al. (2010); Tiede et al. (2020); Jiménez-Rosales et al. (2020); Matsumoto et al. (2020); Dokuchaev and Nazarova (2021); Bauböck et al. (2020); Levis et al. (2022); Wielgus et al. (2022b); Vos et al. (2022); Rosa et al. (2022); Guo et al. (2023); Aimar et al. (2023); Vincent et al. (2024); Rosa et al. (2023); Najafi-Ziyazi et al. (2023); Yfantis et al. (2024b); Levis et al. (2024); Chen et al. (2024a); von Fellenberg et al. (2023); Huang et al. (2024a); Rosa et al. (2024); Chen et al. (2024b); Kocherlakota et al. (2024); Antonopoulou and Nathanail (2024); Yfantis et al. (2024a). This task predominantly employs backward ray tracing, a method that traces geodesics from the observer’s image plane back toward the BH Luminet (1979), and is a foundational technique in most ray tracing and radiative transfer packages Dexter (2016); Moscibrodzka and Gammie (2018); Bronzwaer et al. (2018); Pihajoki et al. (2018); Younsi et al. (2019); Bronzwaer et al. (2020); White (2022); Aimar et al. (2024); Prather et al. (2023); Huang et al. (2024b). However, efficiently simulating hotspot echoes presents challenges. The implementation of a finite grid on the image plane complicates the precise targeting of pointlike sources. While increasing the emission volume may assist, the probability of capturing higher-order images exponentially decreases with each additional image order. Moreover, accurately reconstructing the spatial distribution of emissions within the hotspot is challenging, as only a limited number of geodesics intersect the emission volume. These challenges extend to identifying strongly lensed photons and precisely determining hotspot locations and BH parameters from observational data.
Unlike backward ray tracing, which traces light paths from the observer back to the source, forward ray tracing initiates geodesics at the source and follows them to a distant observer. The main challenge is determining the initial directions of geodesics that reach the observer. One approach is to solve the geodesic equations in integral form using conserved quantities Carter (1968); Bardeen (1973); Dexter and Agol (2009); Lupsasca et al. (2024) (see Eqs. (6) and (7) below) and search for images on the conserved quantity plane. This method has been applied to both localized point sources (e.g., Refs. Cunningham and Bardeen (1973); Vazquez and Esteban (2004)) and extended emitters (e.g., Refs. Beckwith and Done (2005); Gelles et al. (2021)).
In this work, we refine this approach by implementing the recent analytic geodesic solutions from Refs. Gralla and Lupsasca (2020b, a), which provide a complete classification of geodesic motion types and analytic expressions for all the integrals in Eqs. (7). This enables efficient and precise calculations of geodesic directions in all cases. Additionally, we introduce a tailored parametrization scheme for root-finding, facilitating the computation of higher-order images that exponentially converge to the critical curve. As a result, our method is more efficient, accurate, and comprehensive in characterizing images of pointlike sources—especially higher-order images—while remaining applicable to general source and observer locations, improving upon previous works that relied on numerical integration Cunningham and Bardeen (1973); Dolence et al. (2009) or were restricted to specific source locations, such as being infinitely far from the BH in Ref. Vazquez and Esteban (2004).
@@
Additionally, by introducing perturbations to the geodesics Cunningham and Bardeen (1973); Ohanian (1987), we map a finite emission volume onto specific regions of the image plane, enabling the quantification of image shapes and amplification rates. We confirm that higher-order images undergo spatial compression perpendicular to the critical curve due to the photon ring’s instability, while in the parallel direction, these images may either compress or expand depending on the specific location of the emission. We apply our forward ray tracing method to realistic observation scenarios involving highly time-variable hotspots near M87 or Sgr A, conducting analyses to determine the hotspot location and the BH parameters. This approach demonstrates significant promise for elucidating the BH spin with the next-generation EHT (ngEHT) Emami et al. (2023); Johnson et al. (2023); Ayzenberg et al. (2023).
@@
The structure of this paper is as follows: Section II offers an overview of null geodesics within Kerr spacetime and the backward ray tracing method. Section III presents the forward ray tracing method and its application in generating both direct and higher-order images from a point source. In Sec. IV, a perturbative mapping method is presented, which translates emissions from a finite volume to an array of regions on the image plane, allowing for a systematic analysis of image distortion as the order increases. Section V explores the application of forward ray tracing to spacetime tomography and hotspot localization. Section VI summarizes our findings and discusses potential future directions for both theoretical and observational studies. Numerical code KerrP2P Ker (2024) used in this study is available  [^1] KerrP2P: https://github.com/AuroraDysis/KerrP2P. Throughout this study, we adopt a system of geometrized units where the gravitational constant $G$ and the speed of light $c$ are both set to $G=c=1$.
@@ II Null Geodesics in Kerr Spacetime and Backward Ray Tracing
This section provides an overview of null geodesics in Kerr spacetime, focusing on the analytic integral formalism as detailed in Refs. Gralla and Lupsasca (2020b, a). We explore the trajectories of light rays that reach a distant observer, examining their motion and detailing how they are governed by two conserved quantities in Kerr spacetime. These quantities are then mapped to specific coordinates on the image plane. We conclude this section with a brief discussion of the traditional backward ray tracing method used to calculate BH images, paving the way for the discussion on the forward ray tracing method in the following section.
@@ II.1 Null Geodesics in Kerr Spacetime
where the superscripts ${+,-,f,s}$ correspond to the evaluation of $\mathcal{G}_{j}$ at $\theta=\theta_{+}$, $\theta_{-}$, $\theta_{f}$, and $\theta_{s}$, respectively. Ref. Gralla and Lupsasca (2020b) provides the analytic expressions for $\mathcal{G}_{j}$.
@@
where $\gamma$ is the Lyapunov exponent characterizing the instability of bound orbits, as defined in Refs. Cardoso et al. (2009); Gralla and Lupsasca (2020a). Conserved quantities of critical photons correspond to a curve known as the “critical curve” on the $\lambda$-$\eta$ plane, which delineates two distinct regions in the positive-$\eta$ portion of the $\lambda$-$\eta$ plane, each corresponding to a different type of radial motion.
The classification of radial motion leverages the roots of the radial potential $\mathcal{R}(r)$, denoted as ${r_{1},r_{2},r_{3},r_{4}}$, arranged such that $\operatorname{Re}r_{1}\leq\operatorname{Re}r_{2}\leq\operatorname{Re}r_{3}\leq\operatorname{Re}r_{4}$. Detailed analytic expressions for these roots are available in Ref. Gralla and Lupsasca (2020b). The behavior of $r_{4}$ is crucial in determining the nature of radial motion. When $(\lambda,\eta)$ falls inside the critical curve, $r_{4}$ is either complex or real but less than the outer horizon at $r_{h}\equiv M+\sqrt{M^{2}-a^{2}}$. In this scenario, the radial potential $\mathcal{R}(r)$ does not admit zeros outside the horizon, indicating no radial turning points are present, and only photons with $\nu_{r}^{s}=+1$ can escape. We denote the number of radial turning points encountered by $w$, which is zero ($w=0$) in this scenario.
@@
Table 1: Classification of radial motion capable of escaping the BH, delineated by the number $w$ of radial turning points encountered by the photon. CC denotes the critical curve.
@@
with the superscripts $f$, $s$, and $4$ indicating the evaluation of $\mathcal{I}_{i}$ at $r=r_{f}$, $r_{s}$, and $r_{4}$, respectively. Analytic formulas for the antiderivative $\mathcal{I}_{i}(r)$ (for $i=r,\phi,t$) are detailed in Ref. Gralla and Lupsasca (2020b).
@@ II.2 Image Plane
In practice, simulating an image of a BH surrounded by an accretion flow often involves backward ray tracing. This method launches geodesics from a grid on the observer’s image plane, parametrized by $(\alpha,\beta)$, back toward the BH Bardeen (1973); Luminet (1979). The geodesics’ endpoints either fall into the BH or extend to a large Mino time for photons situated outside or on the critical curve. The next step involves calculating the emissivity and absorption coefficients along each geodesic. These values are then integrated back to the observer, factoring in both gravitational and Doppler redshift effects, in a procedure termed covariant radiative transfer Gammie and Leung (2012); Dexter (2016).
@@ III Forward Ray Tracing and Light Echoes
The backward ray tracing method, while effective in many applications, is inefficient for simulating higher-order images from pointlike sources such as hotspots. This inefficiency arises from the challenge of aligning light rays from a finite observer grid precisely with the emission point. Moreover, the photon ring’s inherent instability further complicates this process, as the probability of correctly capturing these light rays decreases exponentially with increasing image order $N$. To address these challenges and facilitate the simulation of hotspot images, we employ forward ray tracing Cunningham and Bardeen (1973); Vazquez and Esteban (2004). Utilizing the analytic geodesic solutions discussed in Sec. II.1 and an optimized parametrization scheme for conserved quantities, our forward ray tracing method effectively captures multiple images from a point source.
@@ III.1 Forward Ray Tracing
This subsection elaborates on the methodology behind our forward ray tracing approach. This approach, initiating at the source point, computes light rays that ultimately converge at the observer, yielding observable images on the plane described in Sec. II.2. Note that our discussion is confined to scenarios where $\eta>0$, thus omitting vortical motion.
@@ III.1.1 Boundary Conditions to Fix Light Rays
Within the refined parameter space, each $(\lambda,\eta,\nu_{r}^{s},\nu_{\theta}^{s})$ combination corresponds to a light ray directed toward a distant observer. Next, we consider the light ray’s final direction. We begin by setting the final radial position to $r_{f}=r_{o}$, which allows us to determine the Mino time $\tau\equiv I_{r}$ through Eq. (16). Subsequently, the final polar angle $\theta_{f}$ is obtained from Eq. (11). This enables us to calculate the number of angular turnings $m$ encountered during propagation
@@
The concluding requirement for a light ray is to satisfy the imaging conditions Cunningham and Bardeen (1973)
@@
Figure 1: Parametrization of conserved quantities of geodesics used in root-finding. The example shown is for a BH with $a/M=0.8$. The blue curve represents the critical curve corresponding to bound orbits, with arrows indicating the direction of increasing $\tilde{r}$.
In practice, to efficiently identify the roots with high precision, we transform the variables $(\lambda,\eta)$ into a new pair of parameters. We begin by introducing $q\equiv\sqrt{\eta}$, ensuring $\lambda$ and $q$ share the same dimensionality. The value of $q$ along the critical curve is denoted as $\tilde{q}\equiv\sqrt{\tilde{\eta}}$. We then assess the deviation of $(\lambda,q)$ from their critical curve counterparts $(\tilde{\lambda},\tilde{q})$ via the parametrization
@@
represents the outward-pointing unit normal vector to the critical curve $(\tilde{\lambda},\tilde{q})$ on the $\lambda$-$q$ plane. The parameter $d$ indicates the distance of $(\lambda,q)$ from the critical curve, with regions outside and inside the curve distinguished by $d>0$ and $d<0$, respectively. This approach allows each point $(\lambda,q)$ to be uniquely mapped to a distance $d$ and the radius $\tilde{r}$ of the corresponding critical curve, as illustrated in Fig. 1. Given that higher-order images asymptotically approach the critical curve, we opt for $\log_{10}|d/M|$ as the variable to accommodate the asymptotic behavior of radial motion. This parametrization scheme effectively alleviates the numerical challenges arising from the exponentially increased sensitivity of $(\theta_{f},\phi_{f})$ on conserved quantities as higher-order images approach the critical curve. The root-finding process is then transformed from the original $\lambda$-$\eta$ plane to the $\tilde{r}$-$\log_{10}|d/M|$ plane, with a total of eight possible combinations of sign variables $(\nu_{r}^{s},\nu_{\theta}^{s},\operatorname{sgn}(d))=(\pm 1,\pm 1,\pm 1)$, which should be solved one by one. For each combination of sign variables, the contours for $\theta_{f}=\theta_{o}$ and $\phi_{f}=\phi_{o}+2k\pi\ (k\in\mathbb{Z})$ are drawn, with their intersections denoting the images. Note that for a Schwarzschild BH, since $\tilde{r}$ is always $3M$, finding a root in the one-dimensional space of $\log_{10}|d/M|$ suffices. Equivalently, this can be done by restricting to the equatorial plane, setting $\eta=0$, and solving for the remaining conserved quantity $\lambda$. See Sec. IV.2 for details.
Our forward ray tracing method is similar to that used in Ref. Vazquez and Esteban (2004), which employs analytic geodesic solutions for computational efficiency and accuracy. However, that study only considers a point source located infinitely far from the BH ($r_{s}\to\infty$), leading to lensed images (referred to as relativistic images therein) originating solely from the region outside the critical curve in the $\lambda$-$\eta$ plane. In this work, we generalize this methodology to arbitrary source positions, utilizing comprehensive analytic geodesic solutions that account for parameter spaces both inside and outside the critical curve. For instance, in Fig. 4, images appear inside the critical curve for $r_{s}=1.7\,M$ and outside for $r_{s}=10\,M$.